OpenStudy (anonymous):

If f(x) = 2x + 8, find f(9). Answer 26 8 19 18 Just say a,b,c, or D. Thank you so much Ill give you a medal for it. (:

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Thank you so much.

OpenStudy (anonymous):

f(9) = 2*9 + 8 = 26 Just replace x with 9...

OpenStudy (saifoo.khan):

f(x) = 2x + 8 f(9) = 2(9) + 8 f(9) = 18 + 8 f(9) = 26